What is the equation y^2-8y+8x=-16
Answer by stanbon(75887)

What is the equation y^2-8y+8x=-16
----
It is the equation of a parabola opening to the left.
Complete the square:
y^2-8y+16 = -8x-16+16
(y-4)^2 = -8x
x = (-1/8)(y-4)^2
=============================
Cheers,
Stan H.